# MsgABD: EveryShipout initializing macros

### v1

@@ Abstract
Message Passing Neural Networks (MPNNs) are a class of Graph Neural Networks (GNNs) that leverage the graph topology to propagate messages across increasingly larger neighborhoods. The message-passing scheme leads to two distinct challenges: over-smoothing and over-squashing. While several algorithms, e.g. DropEdge and its variants – DropNode, DropAgg and DropGNN – have successfully addressed the over-smoothing problem, their impact on over-squashing remains largely unexplored. This represents a critical gap in the literature as failure to mitigate over-squashing would make these methods unsuitable for long-range tasks. In this work, we take the first step towards closing this gap by studying the aforementioned algorithms in the context of over-squashing. We present novel theoretical results that characterize the negative effects of DropEdge on sensitivity between distant nodes, suggesting its unsuitability for long-range tasks. Our findings are easily extended to its variants, allowing us to build a comprehensive understanding of how they affect over-squashing. We evaluate these methods using real-world datasets, demonstrating their detrimental effects. Specifically, we show that while DropEdge-variants improve test-time performance in short-range tasks, they deteriorate performance in long-range ones. Our theory explains these results as follows: random edge-dropping lowers the effective receptive field of GNNs, which although beneficial for short-range tasks, misaligns the models on long-range ones. This forces the models to overfit to short-range artefacts in the training set, resulting in poor generalization. Our conclusions highlight the need to re-evaluate various methods designed for training deep GNNs, with a renewed focus on modelling long-range interactions.
@@ 1 Introduction
Graph-structured data is ubiquitous – it is found in social media platforms, online retail platforms, molecular structures, transportation networks, and even computer systems. Graph neural networks (GNNs) (Scarselli et al., 2009; Li et al., 2016) are powerful neural models developed for modelling graph-structured data, and have found applications in several real-world scenarios (Gao et al., 2018; You et al., 2020c; a; 2022; b; Ying et al., 2018; Zheng et al., 2022; Monti et al., 2017; Zitnik & Leskovec, 2017; Wale & Karypis, 2006). A popular class of GNNs, called message-passing neural networks (MPNNs) (Gilmer et al., 2017), recursively process neighborhood information using message-passing layers. These layers are stacked to allow each node to aggregate information from increasingly larger neighborhoods, akin to how convolutional neural networks (CNNs) learn hierarchical features for images (LeCun et al., 1989). However, unlike in image-based deep learning, where ultra-deep CNN architectures have led to performance breakthroughs (Szegedy et al., 2015; He et al., 2016), shallow GNNs often outperform deeper models on many graph learning tasks (Zhou et al., 2021b). This is because deep GNNs suffer from unique issues like over-smoothing (Oono & Suzuki, 2020) and over-squashing (Alon & Yahav, 2021), which makes training them notoriously difficult.
@@
Over-smoothing refers to the problem of node representations becoming too similar as they are recursively processed. This is undesirable since it limits the GNN from effectively utilizing the information in the input features. The problem has garnered significant attention from the research community, resulting in a suite of algorithms designed to address it Rusch et al. (2023). Amongst these methods are a collection of random edge-dropping algorithms, including DropEdge (Rong et al., 2020) and its variants – DropNode (Feng et al., 2020), DropAgg Jiang et al. (2023) and DropGNN (Papp et al., 2021) – which act as message-passing reducers.
@@
The other issue specific to GNNs is over-squashing. In certain graph structures, neighborhood size grows exponentially with distance from the source (Chen et al., 2018b), causing information to be lost as it passes through graph bottlenecks (Alon & Yahav, 2021). This limits MPNNs’ ability to enable communication between distant nodes, which is crucial for good performance on long-range tasks. To alleviate over-squashing, several graph-rewiring techniques have been proposed, which aim to improve graph connectivity by adding edges in a strategic manner [^1] Sometimes, along with removal of some edges to preserve statistical properties of the original topology. (Karhadkar et al., 2023; Deac et al., 2022; Black et al., 2023; Nguyen et al., 2023; Alon & Yahav, 2021). In contrast, the DropEdge-variants only remove edges, which should, in principle, amplify over-squashing levels.
@@
The empirical evidence in support of methods designed for training deep GNNs has been majorly collected on short-range tasks. That is, it simply suggests that these methods prevent loss of local information, but it remains inconclusive if they facilitate capturing long-range interactions (LRIs). Of course, on long-range tasks, deeper GNNs are useless if they cannot capture LRIs. This is especially a concern for DropEdge-variants since evidence suggests that alleviating over-smoothing with graph rewiring could exacerbate over-squashing (Nguyen et al., 2023; Giraldo et al., 2023).
@@
Theoretical Contributions. In this work, we uncover the effects of random edge-dropping algorithms on over-squashing in MPNNs. By explicitly computing the expected sensitivity of the node representations to the node features (Topping et al., 2022) (inversely related to over-squashing) in a linear GCN Kipf & Welling (2017), we show that these methods provably reduce the effective receptive field of the model. Precisely speaking, the rate at which sensitivity between distant nodes decays is polynomial w.r.t. the dropping probability. Finally, we extend the existing theoretical results on sensitivity in nonlinear MPNNs Black et al. (2023); Di Giovanni et al. (2023); Xu et al. (2018) to the random edge-dropping setting, again showing that these algorithms exacerbate the over-squashing problem.
Experimental Results. We evaluate the DropEdge-variants on node classification tasks using GCN and GAT Veličković et al. (2018) architectures. Specifically, we assess their performance on homophilic datasets – Cora McCallum et al. (2000) and CiteSeer Giles et al. (1998) – which represent short-range tasks, and heterophilic datasets – Chameleon, Squirrel, TwitchDE Rozemberczki et al. (2021) – which correspond to long-range tasks. Our results indicate an increasing trend in test accuracy for homophilic datasets and a declining trend for heterophilic datasets as the dropping probability increases. Accordingly, we hypothesize that edge-dropping algorithms improve performance on short-range tasks, as has been reported earlier, by reducing the receptive field of the GNN and increasing model-dataset alignment, However, for long-range tasks, they decrease the model-dataset alignment, resulting in poor generalization.
@@ 2.1 Graph Neural Networks
Graph neural networks (GNNs) operate on inputs of the form $\left(\mathcal{G},\bm{X}\right)$, where $\mathcal{G}$ encodes the graph topology and $\bm{X}\in^{N\times H^{\left(0\right)}}$ collects the node features [^2] To keep things simple, we will ignore edge features.. Message-passing neural networks (MPNNs) Gilmer et al. (2017) are a special class of GNNs which recursively aggregate information from the 1-hop neighborhood of each node using message-passing layers. An L-layer MPNN is given as
@@
where $\bm{Z}^{\left(0\right)}=\bm{X}$, $\mathsf{Agg}^{\left(\ell\right)}$ denotes the aggregation functions, $\mathsf{Upd}^{\left(\ell\right)}$ the update functions, and $\mathsf{Out}$ the readout function. Since the final representation of node $i$ is a function of the input features of nodes at most L-hops away from it, its receptive field is given by $\mathbb{B}^{\left(L\right)}\left(i\right)\coloneqq\left\{j\in\mathcal{V}:d_{\mathcal{G}}\left(j,i\right)\leq L\right\}$.
For example, a GCN Kipf & Welling (2017) updates node representations as the weighted sum of its neighbors’ representations:
@@
where $\sigma$ is a point-wise nonlinearity, e.g. ReLU, the propagation matrix, $\hat{\bm{A}}$, is a graph shift operator, i.e. $\hat{\bm{A}}_{ij}\neq 0$ if and only if $\left(j\to i\right)\in\mathcal{E}$ or $i=j$, and $\bm{W}^{\left(\ell\right)}\in^{H^{\left(\ell-1\right)}\times H^{\left(\ell\right)}}$ is a weight matrix. The original choice for $\hat{\bm{A}}$ was the symmetrically normalized adjacency matrix $\hat{\bm{A}}^{\text{sym}}\coloneqq\widetilde{\bm{D}}^{-1/2}\widetilde{\bm{A}}\widetilde{\bm{D}}^{-1/2}$ Kipf & Welling (2017), where $\widetilde{\bm{A}}=\bm{A}+\bm{I}_{N}$ and $\widetilde{\bm{D}}=\text{diag}(\widetilde{\bm{A}}\mathbf{1}_{N})$. However, several influential works have also used the asymmetrically normalized adjacency, $\hat{\bm{A}}^{\text{asym}}\coloneqq\widetilde{\bm{D}}^{-1}\widetilde{\bm{A}}$ Hamilton et al. (2017); Schlichtkrull et al. (2017); Li et al. (2018).
2.2 DropEdge
DropEdge is a random data augmentation technique that works by sampling a subgraph of the original input graph in each layer, and uses that for message passing Rong et al. (2020):
$\begin{split}\mathbf{M}^{\left(\ell\right)}&\sim\left\{\text{Bern}\left(1-q\right)\right\}^{N\times N}\\ \widetilde{\bm{A}}^{\left(\ell\right)}&=\mathbf{M}^{\left(\ell\right)}\circ\bm{A}+\bm{I}_{N}\end{split}$ (2.3)
Several variants of DropEdge have also been proposed, forming a family of random edge-dropping algorithms for tackling the over-smoothing problem. For example, DropNode (Feng et al., 2020) independently samples nodes and sets their representations to 00, followed by rescaling to make the feature matrix unbiased. This is equivalent to setting the corresponding columns of the propagation matrix to 00. In a similar vein, DropAgg (Jiang et al., 2023) samples nodes that don’t aggregate messages from their neighbors. This is equivalent to dropping the corresponding rows of the adjacency matrix. Combining these two approaches, DropGNN (Papp et al., 2021) samples nodes which neither propagate nor aggregate messages in a given message-passing step. These algorithms alleviate over-smoothing by reducing the number of messages being propagated in the graph, thereby slowing down the convergence of node representations.
@@ 2.3 Over-squashing
Over-squashing refers to the problem of information from exponentially growing neighborhoods Chen et al. (2018a) being squashed into finite-sized node representations Alon & Yahav (2021). This results in a loss of information as it is propagated over long distances, disallowing MPNNs from capturing long-range interactions (LRIs) and limiting their applications to short-range tasks. Topping et al. (2022) formally characterized over-squashing in terms of the Jacobian of the node-level representations w.r.t. the input features: $\|\partial\bm{z}^{\left(L\right)}_{i}/\partial\bm{x}_{j}\|_{1}$. Accordingly, over-squashing can be understood as low sensitivity between distant nodes, i.e. small perturbations in a node’s features don’t effect other distant nodes’ representations.
Several works have linked over-squashing in an MPNN with topological properties like Cheeger’s constant Giraldo et al. (2023); Karhadkar et al. (2023), curvature of edges Topping et al. (2022); Nguyen et al. (2023); Liu et al. (2023), effective resistance between nodes Black et al. (2023); Arnaiz-Rodríguez et al. (2022) and the expected commute time between them Di Giovanni et al. (2023); Giovanni et al. (2024). These results have inspired the design of several graph rewiring techniques that strategically add edges to improve the connectivity in the graph, thereby alleviating over-squashing.
@@ 3 Sensitivity Analysis
3.1 Linear GCNs
We start our analysis with linear GCNs, and treat more general MPNN architectures in the following subsection. In this model, the final node representations can be summarised as
@@ 3.1 Linear GCNs
See Section B.1 for a proof.
Other Variants. We will similarly derive the expected propagation matrix for other random edge-dropping algorithms. First off, DropNode (Feng et al., 2020) samples nodes and drops corresponding columns from the aggregation matrix directly, followed by rescaling of its entries:
$\begin{split}\mathbb{E}_{\mathsf{DN}}\left[\frac{1}{1-q}\hat{\bm{A}}\right]=\frac{1}{1-q}\times\left(1-q\right)\hat{\bm{A}}=\hat{\bm{A}}\end{split}$ (3.4)
Nodes sampled by DropAgg Jiang et al. (2023) don’t aggregate messages. Therefore, if $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, then the expected propagation matrix is given by
$\begin{split}\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]&=q+\frac{1-q}{d_{i}+1}=\frac{1+d_{i}q}{d_{i}+1}>\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ii}\right]\\ \mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ij}\right]&=\frac{1}{d_{i}}\left(1-\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]\right)<\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ij}\right]\end{split}$ (3.5)
Finally, DropGNN (Papp et al., 2021) samples nodes which neither propagate nor aggregate messages. From any node’s perspective, if it is not sampled, then its aggregation weights are computed as for DropEdge:
$\begin{split}\mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ii}\right]&=q+\left(1-q\right)\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ii}\right]=q+\frac{1-q^{d_{i}+1}}{d_{i}+1}>\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]\\ \mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ij}\right]&=\frac{1}{d_{i}}\left(1-\mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ii}\right]\right)<\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ij}\right]\end{split}$ (3.6)
@@
where the right-hand sides of the two inequalities are the corresponding entries in the propagation matrix of a NoDrop model. Equation 3.3 to Equation 3.7 together imply the following result:
@@
In other words, DropEdge-variants prevent a 1-layer GCN from fully utilizing neighborhood information when learning node representations. Ignoring the graph topology this way makes the model resemble an MLP, limiting its expressiveness and hindering its ability to model graph-data.
Figure 1: Entries of $\dot{\bm{P}}^{6}$, averaged over molecular graphs sampled from the Proteins dataset. Left: Sensitivity between nodes decays at a polynomial rate w.r.t. their distance. Middle: Similarly, it decays at a polynomial rate w.r.t. the DropEdge probability. Right: It decays more quickly with DropAgg than with DropEdge, and even more rapidly with DropGNN.
L-layer Linear GCNs. Unfortunately, we cannot draw similar conclusions in L-layer networks, for nodes at arbitrary distances. To see this, view $\dot{\bm{P}}$ as the transition matrix of a non-uniform random walk. This walk has higher self-transition ($i=j$) probabilities than in a uniform augmented random walk ($\bm{P}=\hat{\bm{A}}^{\text{asym}}$, $q=0$), but lower inter-node ($i\neq j$) transition probabilities. Note that $\dot{\bm{P}}^{L}$ and $\bm{P}^{L}$ store the L-step transition probabilities in the corresponding walks. Then, since the paths connecting the nodes $i\in\mathcal{V}$ and $j\in\mathbb{B}^{\left(L-1\right)}\left(i\right)$ may involve self-loops, $(\dot{\bm{P}}^{L})_{ij}$ may be lower or higher than $(\bm{P}^{L})_{ij}$. Therefore, we cannot conclude how sensitivity between nodes separated by at most $L-1$ hops changes. For nodes L-hops away, however, we can show that DropEdge always decreases the corresponding entry in $\dot{\bm{P}}^{L}$, reducing the effective reachability of GCNs. Using Equation 3.5 and Equation 3.6, we can show the same for DropAgg and DropGNN, respectively.
@@
In an L-layer linear GCN with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, using DropEdge, DropAgg or DropGNN decreases the sensitivity of a node $i\in\mathcal{V}$ to another node $j\in\mathbb{S}^{\left(L\right)}\left(i\right)$, thereby reducing its effective receptive field. Moreover, the sensitivity monotonically decreases as the dropping probability is increased.
@@
Nodes at Arbitrary Distances. Although no general statement could be made about the change in sensitivity between nodes up to $L-1$ hops away, we can analyze such pairs empirically. We sampled 100 molecular graphs from the Proteins dataset (Dobson & Doig, 2003), binned the node-pairs in each graph by the shortest distance between them, and then plotted the average of the corresponding entries in $\dot{\bm{P}}^{L}$, $L=6$.
The results are shown in Figure 1. In the left subfigure, we observe that the sensitivity between two nodes decays at a polynomial rate with increasing distance between them. Moreover, DropEdge increases the expected sensitivity between nodes close to each other (0-hop and 1-hop neighbors) in the original topology, but reduces it between nodes farther off. In the middle subfigure, we show how the average sensitivity at different distances changes with the DropEdge probability. Specifically, we observe that the decay in sensitivity to nodes at large distances is polynomial in the DropEdge probability, which suggests that the algorithm would not be suitable for capturing LRIs. Similar conclusions can be made with the symmetrically normalized propagation matrix (see Section D.1). Finally, in the right subfigure, we compare the DropEdge-variants by plotting the sensitivity at $d_{\mathcal{G}}\left(i,j\right)=L=6$. Although the analytical form of all the expected propagation matrices are available to us, we approximate them using Monte-Carlo sampling. As expected from Equation 3.5, we observe that DropAgg not only decreases the sensitivity of node representations to distant nodes, but does it to a greater extent than DropEdge. Similarly, Equation 3.6 suggested that DropGNN could be even more harmful than DropAgg, and this is validated by the empirical results. In fact, for $q>0.7$, nodes 6-hops away are mostly insensitive to each other.
3.2 Nonlinear MPNNs
ReLU GCNs. Xu et al. (2018) considered the case of ReLU nonlinearity, so that the update rule is $\bm{Z}^{\left(\ell\right)}=\mathsf{ReLU}(\hat{\bm{A}}\bm{Z}^{\left(\ell-1\right)}\bm{W}^{\left(\ell\right)})$. Additionally, it makes the simplifying assumption that each path in the computational graph is active with a fixed probability, $\rho$ (Kawaguchi, 2016, Assumption A1p-m). Accordingly, the sensitivity (in expectation) between any two nodes is given as
$\left\|\mathbb{E}_{\mathsf{ReLU}}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}=\left[\rho\left\|\prod_{\ell=1}^{L}\bm{W}^{\left(\ell\right)}\right\|_{1}\right]\left(\hat{\bm{A}}^{L}\right)_{ij}=\zeta_{1}^{\left(L\right)}\left(\hat{\bm{A}}^{L}\right)_{ij}$ (3.8)
where $\zeta_{1}^{\left(L\right)}$ is independent of the choice of nodes $i,j\in\mathcal{V}$. Taking an expectation w.r.t. the random edge masks, we get
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\mathbb{E}_{\mathsf{ReLU}}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}\right]$ $=\zeta_{1}^{\left(L\right)}\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left(\prod_{\ell=1}^{L}\hat{\bm{A}}^{\left(\ell\right)}\right)_{ij}\right]$ (3.9)
$=\zeta_{1}^{\left(L\right)}\left(\mathbb{E}\left[\hat{\bm{A}}\right]^{L}\right)_{ij}$ (3.10)
Using Theorem 3.1, we conclude that in a ReLU-GCN, DropEdge, DropAgg and DropGNN will reduce the expected sensitivity between nodes L-hops away. Empirical observations in Figure 1 and Figure 4 suggest that we may expect an increase in sensitivity to neighboring nodes, but a significant decrease in sensitivity to those farther away.
Source-only Message Functions. Black et al. (2023, Lemma 3.2) considers MPNNs with aggregation functions of the form

### v2

# Jasraj Singh Independent Researcher &Keyue Jiang

@@ Abstract
Message Passing Neural Networks (MPNNs) are a class of Graph Neural Networks (GNNs) that propagate information across the graph via local neighborhoods. The scheme gives rise to two key challenges: over-smoothing and over-squashing. While several Dropout-style algorithms, such as DropEdge and DropMessage, have successfully addressed over-smoothing, their impact on over-squashing remains largely unexplored. This represents a critical gap in the literature, as failure to mitigate over-squashing would make these methods unsuitable for long-range tasks – the intended use case of deep MPNNs. In this work, we study the aforementioned algorithms, and closely related edge-dropping algorithms – DropNode, DropAgg and DropGNN – in the context of over-squashing. We present theoretical results showing that DropEdge-variants reduce sensitivity between distant nodes, limiting their suitability for long-range tasks. To address this, we introduce DropSens, a sensitivity-aware variant of DropEdge that explicitly controls the proportion of information lost due to edge-dropping, thereby increasing sensitivity to distant nodes despite dropping the same number of edges. Our experiments on long-range synthetic and real-world datasets confirm the predicted limitations of existing edge-dropping and feature-dropping methods. Moreover, DropSens consistently outperforms graph rewiring techniques designed to mitigate over-squashing, suggesting that simple, targeted modifications can substantially improve a model’s ability to capture long-range interactions. Our conclusions highlight the need to re-evaluate and re-design existing methods for training deep GNNs, with a renewed focus on modelling long-range interactions.
@@ 1 Introduction
Graph neural networks (GNNs) [71, 51] are powerful neural models developed for modelling graph-structured data, and have found applications in several real-world scenarios [29, 91, 89, 92, 90, 88, 94, 60, 99, 82]. A popular class of GNNs, called message-passing neural networks (MPNNs) [32], recursively process neighborhood information using message-passing layers. These layers are stacked to allow each node to aggregate information from increasingly larger neighborhoods, akin to how convolutional neural networks (CNNs) learn hierarchical features for images [48]. However, unlike in image-based deep learning, where ultra-deep CNN architectures have led to performance breakthroughs [78, 38], shallow GNNs often outperform deeper models on many graph learning tasks [97]. This is because deep GNNs suffer from unique issues like over-smoothing [64] and over-squashing [4], which makes training them notoriously difficult.
@@
Over-smoothing refers to the problem of node representations becoming too similar as they are recursively processed. This is undesirable since it limits the GNN from effectively utilizing the information in the input features. The problem has garnered significant attention from the research community, resulting in a suite of algorithms designed to address it [70] (see Section A.1 for an overview of representative methods). Amongst these methods are a collection of random edge-dropping algorithms, including DropEdge [68], DropNode [23], DropAgg [43] and DropGNN [65] – which we will collectively refer to as DropEdge-variants – which act as message-passing reducers. In addition, we have DropMessage [21], which performs Dropout [77] on the message matrices, instead of the feature matrices; we will collectively refer to these two methods as Dropout-variants since they are applied along the feature dimensions.
@@
The other issue specific to GNNs is over-squashing. In certain graph structures, neighborhood size grows exponentially with distance from the source [12], causing information to be lost as it passes through graph bottlenecks [4]. This limits MPNNs’ ability to enable communication between distant nodes, which is crucial for good performance on long-range tasks. To alleviate over-squashing, several graph-rewiring techniques have been proposed, which aim to improve graph connectivity by adding edges in a strategic manner [44, 16, 8, 63, 4] (see Section A.3 for an overview of representative methods). [^1] Sometimes, along with removal of some edges to preserve statistical properties of the original topology. In contrast, the DropEdge-variants only remove edges, which should, in principle, amplify over-squashing levels. The same can be intuitively argued about Dropout-variants.
@@
Empirical evidence in support of methods designed for training deep GNNs has been majorly collected on short-range tasks (see Section A.2 for a detailed discussion). That is, it simply suggests that these methods prevent loss of local information, but it remains inconclusive if they facilitate capturing long-range interactions (LRIs). Of course, on long-range tasks, deep GNNs are useless if they cannot capture LRIs. This is especially a concern for DropEdge-variants since evidence suggests that alleviating over-smoothing with graph rewiring could exacerbate over-squashing [63, 34].
@@
Contributions. In this work, we precisely characterize the effects of random edge-dropping algorithms on over-squashing in MPNNs. By explicitly computing the expected sensitivity of the node representations to the node features [79] (inversely related to over-squashing) in a linear Graph Convolutional Network (GCN) [47], we show that these methods provably reduce the effective receptive field of the model. Precisely speaking, the rate at which sensitivity between nodes decays is exponential w.r.t. the distance between them. We also extend the existing theoretical results on sensitivity in nonlinear MPNNs [8, 18, 84] to the random edge-dropping setting, concluding that these algorithms exacerbate the over-squashing problem. We use our analysis of GCNs to design a sensitivity-aware DropEdge-variant, named DropSens, that enjoys the representational expressivity of DropEdge without suffering from over-squashing, thereby demonstrating how algorithms can be readily adapted for long-range tasks.
We evaluated the DropEdge- and Dropout-variants on long-range datasets using GCN, Graph Isomorphism Network (GIN) [85] and Graph Attention Network (GAT) [81] architectures. Specifically, we follow the setup in Giovanni et al. [33] with the SyntheticZINC dataset, in Topping et al. [79] with real-world homophilic (corresponding to short-range tasks) and heterophilic (long-range tasks) node classification datasets, and in Black et al. [8], Karhadkar et al. [44] with graph classification datasets. Our results indicate that while the random dropping methods improve model performance in short-range tasks, they are often ineffective, and sometimes even detrimental, to long-range task performance. Finally, we present results for DropSens, which outperforms state-of-the-art graph rewiring methods aimed at addressing over-squashing at node classification and graph-classification tasks. These findings point to the importance of re-evaluating the methods used to train deep GNNs, especially in terms of how well they capture LRIs.
@@ 2.1 Graph Neural Networks
Graph Neural Networks (GNNs) operate on inputs of the form $\left(\mathcal{G},\bm{X}\right)$, where $\mathcal{G}$ encodes the graph topology and $\bm{X}\in{}^{N\times H^{\left(0\right)}}$ collects the node features. [^2] To keep things simple, we will ignore edge features. Message-Passing Neural Networks (MPNNs) [32] are a special class of GNNs which recursively aggregate information from the 1-hop neighborhood of each node using message-passing layers. An L-layer MPNN is given as
@@
where $\bm{Z}^{\left(0\right)}=\bm{X}$, $\mathsf{Agg}^{\left(\ell\right)}$ denotes the aggregation functions, $\mathsf{Upd}^{\left(\ell\right)}$ the update functions, and $\mathsf{Out}$ the readout function. Since $\bm{z}_{i}^{\left(L\right)}$ is a function of the input features of nodes at most L-hops away from it, its receptive field is given by $\mathbb{B}^{\left(L\right)}\left(i\right)\coloneqq\left\{j\in\mathcal{V}:d_{\mathcal{G}}\left(j,i\right)\leq L\right\}$.
For example, a GCN [47] updates node representations as the weighted sum of its neighbors’ representations:
@@
where $\sigma$ is a point-wise nonlinearity, e.g. ReLU, the propagation matrix, $\hat{\bm{A}}$, is a graph shift operator, i.e. $\hat{\bm{A}}_{ij}\neq 0$ if and only if $\left(j\to i\right)\in\mathcal{E}$ or $i=j$, and $\bm{W}^{\left(\ell\right)}\in{}^{H^{\left(\ell-1\right)}\times H^{\left(\ell\right)}}$ is a weight matrix. The original choice for $\hat{\bm{A}}$ was the symmetrically normalized adjacency matrix $\hat{\bm{A}}^{\text{sym}}\coloneqq\widetilde{\bm{D}}^{-1/2}\widetilde{\bm{A}}\widetilde{\bm{D}}^{-1/2}$ [47], where $\widetilde{\bm{A}}=\bm{A}+\bm{I}_{N}$ and $\widetilde{\bm{D}}=\text{diag}(\widetilde{\bm{A}}\mathbf{1}_{N})$. However, several influential works have also used the asymmetrically normalized adjacency, $\hat{\bm{A}}^{\text{asym}}\coloneqq\widetilde{\bm{D}}^{-1}\widetilde{\bm{A}}$ [36, 72, 50].
2.2 DropEdge-variants
DropEdge [68] is a random data augmentation technique that works by sampling a subgraph of the input graph in each layer, followed by the addition of self-loops, and uses that for message passing. Several variants of DropEdge have also been proposed, forming a family of random edge-dropping algorithms for tackling the over-smoothing problem. For example, DropNode [23] independently samples nodes and sets their features to $\mathbf{0}$, followed by rescaling to make the feature matrix unbiased. This is equivalent to setting the corresponding columns of the propagation matrix to $\mathbf{0}$. In a similar vein, DropAgg [43] samples nodes that don’t aggregate messages from their neighbors. This is equivalent to dropping the corresponding rows of the adjacency matrix. Combining these two approaches, DropGNN [65] samples nodes which neither propagate nor aggregate messages in a given layer. These algorithms alleviate over-smoothing by reducing the number of messages being propagated in the graph, thereby slowing down the convergence of node representations.
2.3 Dropout-variants
Dropout is a stochastic regularization technique which reduces over-fitting by randomly dropping features before each layer. It has been successful with various architectures, like CNNs [77] and transformers [80], and has also found applications in GNN training. DropMessage [21] is a variant of Dropout designed specifically for message-passing schemes – it acts directly on the messages over each edge, instead of the node representations. This reduces the induced variance in the messages compared to Dropout, DropEdge and DropNode, while at the same time making the method more effective at alleviating over-smoothing and enabling the training of deep GNNs.
@@ 2.4 Over-squashing
Over-squashing refers to the problem of information from exponentially growing neighborhoods [11] being squashed into finite-sized node representations [4]. Topping et al. [79] formally characterized over-squashing in terms of the Jacobian of the node-level representations w.r.t. the input features: $\|\partial\bm{z}^{\left(L\right)}_{i}/\partial\bm{x}_{j}\|_{1}$. Accordingly, over-squashing can be understood as low sensitivity between distant nodes, i.e. small perturbations in a node’s features don’t effect other distant nodes’ representations.
See Appendix A for an extensive discussion of related works addressing the problems of over-smoothing and over-squashing, and a unified treatment of the two.
@@ 3 Sensitivity Analysis
Here, we present our analysis for linear GCNs, and treat more general nonlinear MPNN architectures in Section C.1. In this model, the final node representations can be summarised as
@@
See Section B.1 for a proof, and a similar treatment of DropNode, DropAgg and DropGNN.
@@
where the right-hand sides of the two inequalities are the corresponding entries in the propagation matrix of a NoDrop model. s 3.3, 3.4, B.16 and B.17 together imply the following result:
@@
(a) Entries of $\dot{\bm{P}}^{6}$ decay at exponential rate w.r.t. distance between nodes, and polynomial rate w.r.t. to the DropEdge probability.
(b) MC-approximation of influence distribution in ReLU-GCNs.
Figure 1: Empirical sensitivity analysis using the Cora dataset.
L-layer Linear GCNs. Unfortunately, we cannot draw similar conclusions in L-layer networks, for nodes at arbitrary distances. To see this, view $\dot{\bm{P}}$ as the transition matrix of a non-uniform random walk. This walk has higher self-transition ($i=j$) probabilities than in a uniform augmented random walk ($\bm{P}=\hat{\bm{A}}^{\text{asym}}$, $q=0$), but lower inter-node ($i\neq j$) transition probabilities. Note that $\dot{\bm{P}}^{L}$ and $\bm{P}^{L}$ store the L-step transition probabilities in the corresponding walks. Then, since the paths connecting the nodes $i\in\mathcal{V}$ and $j\in\mathbb{B}^{\left(L-1\right)}\left(i\right)$ may involve self-loops, $(\dot{\bm{P}}^{L})_{ij}$ may be lower or higher than $(\bm{P}^{L})_{ij}$. Therefore, we cannot conclude how sensitivity between nodes separated by at most $L-1$ hops changes. For nodes L-hops away, however, we can show that DropEdge always decreases the corresponding entry in $\dot{\bm{P}}^{L}$, reducing the effective reachability of GCNs. Using s B.16 and B.17, we can show the same for DropAgg and DropGNN, respectively.
@@
In an L-layer linear GCN with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, using DropEdge, DropAgg or DropGNN decreases the sensitivity of a node $i\in\mathcal{V}$ to another node $j\in\mathbb{S}^{\left(L\right)}\left(i\right)$, thereby reducing its effective receptive field. Moreover, the sensitivity decreases with increasing dropping probability.
@@
Nodes at Arbitrary Distances. Although no general statement could be made about the change in sensitivity between nodes up to $L-1$ hops away, we can analyze such pairs empirically. We compute the L-hop transition matrix $\dot{\bm{P}}^{L}$ – proportional to expected sensitivity in linear GCNs under DropEdge – for the Cora dataset, and average the entries after binning node pairs by the shortest distance between them. The results are shown in 1(a). In the left subfigure, we observe that the expected sensitivity decays at an exponential rate with increasing distance between the corresponding nodes. In the middle subfigure, we observe that DropEdge increases the expected sensitivity between nodes close to each other (0-hop and 1-hop neighbors) in the original topology, but reduces it between nodes farther off. Similar conclusions can be made with the symmetrically normalized propagation matrix (see Section D.1). Note that the over-squashing effects of DropAgg and DropGNN would, in theory, be even more severe, as suggested by s B.16 and B.17.
Nonlinear MPNNs. While linear networks are useful in simplifying the theoretical analysis, they are often not practical. In Section C.1, we treat the upper bounds on sensitivity established in previous works, and extend them to the DropEdge-variants. Even still, although theoretical bounds offer valuable guarantees, they can be arbitrarily loose in the absence of error quantification, making their practical relevance unclear. To reliably conclude the empirical behaviour of DropEdge- and Dropout-variants, we turn to Monte Carlo simulations with ReLU-GCNs; see Section D.3 for a description of the experiment setup. 1(b) compares the influence of the source nodes [84] at different distances using a dropout probability of $0.5$. We observe that while the effect of DropNode on the sensitivity profile – as compared to the baseline NoDrop – is relatively insignificant, models using DropEdge, DropAgg and DropGNN have remarkably lower sensitivity to distant nodes, as predicted by our theory.
4 Sensitivity-Aware DropEdge
Lemma 3.1 tells us that DropEdge decreases the weight of cross-edges, $\left(j\to i\right)$, in the expected propagation matrix, i.e. the strength of message passing over these edges decreases. The fraction of information preserved over a cross-edge is dependent only on the dropping probability and the target node’s in-degree, $d_{i}$. We can directly control this quantity using a per-edge dropping probability, $q_{i}$, dependent only on the receiving node’s in-degree: